Based on the Rational Root Theorem, which of the following is NOT possible roots for the function: 2x- 9x2 + 13x - 6

a. -3/2

b. 6

c. 1/2

d. -9


1
Expert's answer
2021-04-29T15:49:17-0400

"2x^3- 9x^2 + 13x - 6 \\ p= ± 1, ± 2, ±3, ±6 \\ q= ± 1, ± 2 \\ Possible \ rational \ roots \ are: \ x=p/q=±\{1,2,3,6,\frac{1}{2},\frac{3}{2} \} \\ Answer: \ d. \ x=-9"